A GRADUATED CYLINDERCONTAINS 123mLOF WATER. A 45.0g PIECE OF SILVER (d=10.49g/L) IS ADDED. WHAT IS THE NEW WATER LEVELOF THE GRADUATED CYLINDER, IN mLOF THE CYLINDER? D=m/V
Answer #1

We know that, Density = mass/volume

or d=m/v where mass m is in grams, volume v is in litres, thus units of density are g/L

Given that density of silver = 10.49g/L and its mass = 45g

therefore, volume v =m/d = 45/10.49 = 4.289 Litres \approx 4.29 Litres

as 1L = 1000ml

4.29L = 4290 ml

Earlier, the graduated cylinder had 123ml of water. On addition of the silver piece, new volume will be 123+4290= 4413ml